OpenStudy (anonymous):

PLS tell me if I did this correctly? Give the value of x where the function f(x) ((sqrt(x)-1))/(x-1) has an infinite discontinuity? I rationalized and landed up with 1/(sqrtx+1) which means no infinite disc.

myininaya (myininaya):

\[\frac{\sqrt{x}-1}{x-1}*\frac{\sqrt{x}+1}{\sqrt{x}+1}=\frac{x-1}{(x-1)(\sqrt{x}+1)}=\frac{1}{\sqrt{x}+1}\] you are right there are none

OpenStudy (anonymous):

sweet! thank you! Myin!

myininaya (myininaya):

there is a hole at x=1 but not an infinite discontinuity

OpenStudy (anonymous):

yup, I recognize that now since you guys showed me how it works.